# MaxPoolBERT: Enhancing BERT Classification via Layer- and Token-Wise Aggregation

### v1

@@ 4 Experiments
In order to evaluate each previously presented modification of the BERT architecture for sequence classification, we fine-tune each model on different classification tasks and compare the results.
@@ 4.1 Datasets
  • MRPC (Microsoft Research Paraphrase Corpus (Dolan and Brockett, 2005)): 5,800 sentence pairs from news source, annotated for paraphrase identification (equivalent, not equivalent).
  • QNLI (Question NLI): an NLI dataset derived from the Stanford Question Answering Dataset (SQuAD) (Rajpurkar et al., 2016) containing question paragraph pairs. The task is to predict if the question is answered by the given paragraph (entailment, no entailment).
  • MNLI (Multi-Genre NLI (Williams et al., 2018)): includes 433,000 sentence pairs, annotated with three different indicators for entailment: neutral, contradiction and entailment.
@@
  • STS-B (Semantic Textual Similarity Benchmark (Cer et al., 2017)): 8,630 sentence pairs annotated with a textual similarity score ([0,5]).
  • RTE (Recognizing Textual Entailment (Dagan et al., 2006)): 5,770 sentence pairs annotated for entailment recognition (entailment, no entailment).
  • QQP (Quora Question Pairs): 795,000 pairs of questions from Quora, annotated for semantical similarity (duplicate, no duplicate).
  • WNLI (Winograd NLI (Levesque et al., 2012)): 852 sentence pairs annotated for textual entailment (entailment, no entailment).
4.2 Experiment Details
All experiments were run on a single NVIDIA A100 GPU. We used the Huggingface transformers and dataset libraries [^2] https://huggingface.co/ to implement and train all of our models. Each model was fine-tuned three times for four epochs, with different three different random seeds and report the mean of all runs. All hyperparameters are listed in Table 1. The reported results were run on the validation sets of all GLUE tasks.
@@ 5.1 Performance Across GLUE Tasks
The $\text{Max}_{\text{CLS}}$ variant, which applies max-pooling over the [CLS] token representations from the last $k$ layers, results in marginal to no improvement for most tasks. Notably, this variant achieves the best performance among all variants on QNLI and STS-B, suggesting that layer-wise max pooling can be beneficial for certain task types. Both tasks incorporate semantic matching between two texts, thus both require nuanced understanding of sentence meaning.
The MHA variant introduces an additional MHA layer, allowing the final-layer [CLS] token to attend to the full sequence before classification. This variant consistently improves upon the baseline BERT model, indicating that this extra attention step, effectively enhances the model’s ability to integrate global context. The biggest improvement is observed on the WNLI dataset, which has the fewest training examples in the GLUE benchmark (634 training examples), suggesting that the added attention is particularly helpful in low-resource settings.
@@ 6 Conclusion
We introduced MaxPoolBERT, a lightweight yet effective refinement of BERT’s classification pipeline that improves the representational quality of the [CLS] token. Our method leverages max-pooling across layers and tokens, and introduces a multi-head attention layer that allows the [CLS] token to re-aggregate contextual information before classification. These modifications require no changes to pre-training and add minimal overhead to fine-tuning.
Empirical results on the GLUE benchmark demonstrate that MaxPoolBERT outperforms standard BERT-base across most tasks, with especially strong improvements in low-resource settings. This suggests that BERT’s native use of the final-layer [CLS] embedding underutilizes available information and that small architectural additions can enhance generalization without sacrificing efficiency.
@@
  • Model size and generalization. Our work focuses on BERT-base. It remains to be tested whether these findings generalize to larger models like BERT-large or to other transformer variants.
@@
In the future we aim to further investigate how to optimize the fine-tuning of small BERT models. While larger models often yield better performance, smaller models are crucial in real-time or resource-contrained environments. The BERT training paradigm following pre-training and fine-tuning has been predominant for several years and is widely used, so it is important to study whether further improvements can be made through small changes to this learning paradigm.
Ethics Statement
This work does not involve the collection or use of sensitive personal data, nor does it generate or process harmful content. We use only publicly available datasets (GLUE benchmark) and follow standard evaluation protocols. Our modifications are intended to improve model efficiency and robustness, particularly in low-resource settings. No human subjects were involved in this research.
@@ A.1 Ablations
We conduct ablation studies to test different modifications of out architectures. We describe all experiments and report their results in the following.
@@ A.1.1 Choice of $k$
A.1.2 Pooling.
Several studies indicate, that max-pooling seems to be a stable choice to aggregate information into a single sentence representation. To underline these results, we also experiment with mean-pooling. We apply an element-wise mean-pooling operation
on every vector of our $k$ chosen layers (defined analogously as the max operation in Section 3.1). Here we also choose $k=3$. We report the experimental results on the GLUE validation set in Table 5. Overall, mean pooling performs similarly well as max pooling, on RTE it even achieves higher performance than max-pooling. In the end, max-pooling seems to be a better choice as max-pooling as it works better for most GLUE datasets.

### v2

@@ 4 Experiments
In order to evaluate each previously presented modification of the BERT architecture for sequence classification, we fine-tune each model on different classification tasks of the GLUE benchmark and compare the results. As a baseline, we use a standard BERT base model (Devlin et al., 2019). In addition, we assess the generalizability of our approach by applying it to a BERT variant, namely RoBERTa base (Liu et al., 2019).
@@ 4.1 Datasets
  • MRPC (Microsoft Research Paraphrase Corpus (Dolan & Brockett, 2005)): 5,800 sentence pairs from news source, annotated for paraphrase identification (equivalent or not equivalent).
  • QNLI (Question NLI): an NLI dataset derived from the Stanford Question Answering Dataset (SQuAD) (Rajpurkar et al., 2016) containing question paragraph pairs. The task is to predict if the question is answered by the given paragraph (entailment or no entailment).
  • MNLI (Multi-Genre NLI (Williams et al., 2018)): includes 433,000 sentence pairs, annotated with three different indicators for entailment (neutral, contradiction or entailment). MNLI includes both matched (in-domain) and mismatched (cross-domain) sections.
@@
  • STS-B (Semantic Textual Similarity Benchmark (Cer et al., 2017)): 8,630 sentence pairs annotated with a textual similarity score (from zero to five).
  • RTE (Recognizing Textual Entailment (Dagan et al., 2006)): 5,770 sentence pairs annotated for entailment recognition (entailment or no entailment).
  • QQP (Quora Question Pairs): 795,000 pairs of questions from Quora, annotated for semantical similarity (duplicate or no duplicate).
  • WNLI (Winograd NLI Levesque et al., 2012): 852 sentence pairs annotated for textual entailment (entailment or no entailment).
4.2 Experimental Details
All experiments were run on a single NVIDIA A100 GPU. We used the Huggingface transformers and dataset libraries [^2] https://huggingface.co/ to implement and train all of our models. Each model was fine-tuned three times with three different random seeds for four epochs. We report the mean of all runs and use the validation sets of all GLUE tasks for evaluation. Experimenting with different values for $k$ (the number of the considered layers), we found that $k=3$ works best (see Appendix A.1.1). All others hyperparameters are listed in Table 1.
Model CoLA MRPC QNLI MNLI SST-2 STS-B RTE QQP WNLI
MCC Acc. F1 Acc. m $m$ Acc. Sp. Acc. Acc. F1 Acc.
[…]
BERT base 53.59 82.43 87.49 90.96 84.27 84.57 92.55 88.47 63.42 90.65 87.40 49.77
MHA 55.88 85.38 89.51 90.49 84.37 84.67 92.32 88.04 64.98 90.67 87.45 55.4
$\text{Max}_{\text{Seq}}$+MHA 55.35 85.95 89.78 90.73 83.82 84.24 92.74 88.22 66.06 90.59 87.32 55.4
$\text{Mean}_{\text{Seq}}$+MHA 55.10 85.62 89.66 90.86 83.78 84.2 92.51 87.91 66.67 90.68 87.41 54.46
$\Delta$ 2.29 3.52 2.29 0.19 0.24 0.1 0.19 0.5 3.25 0.03 0.05 5.63
Table 2: Our proposed variants improve the performance over BERT base on GLUE validation tasks (average of 3 seeds). The size of the training data set is highlighted in gray. We report Matthews correlation coefficient (MCC) for CoLA, accuracies for matched (m) and mismatched results ($m$) for MNLI, and Spearman correlation (Sp.) for STS-B. Below we report the improvement from the best performing variant over the baseline as $\Delta$.
@@ 5.1 Performance Across GLUE Tasks
Model GLUE avg.
BERT Base 79.63
$\text{Max}_{\text{CLS}}$ 80.02
MHA 80.76
$\text{Max}_{\text{Seq}}$+MHA 80.85
$\text{Mean}_{\text{Seq}}$+MHA 80.75
$\Delta$ 1.25
Table 4: Average performance across all GLUE tasks. MaxPoolBERT shows a consistent gain over BERT base.
The $\text{Max}_{\text{CLS}}$ variant, which applies max-pooling over the [CLS] token representations from the last $k$ layers, results in marginal to no improvement for most tasks. Notably, this variant achieves the best performance among all variants on QNLI and STS-B, suggesting that layer-wise max-pooling can be beneficial for certain task types. Both tasks incorporate semantic matching between two texts, thus both require nuanced understanding of sentence meaning.
The MHA variant introduces an additional MHA layer, allowing the final-layer [CLS] token to attend to the full sequence before classification. This variant consistently improves upon the baseline BERT model, indicating that this extra attention step, effectively enhances the model’s ability to integrate global context. The biggest improvement is observed on the WNLI dataset, which has the fewest training examples in the GLUE benchmark (634 training examples in total), suggesting that the added attention is particularly helpful in low-resource settings.
@@
Overall, $\text{Mean}_{\text{Seq}}$+MHA performs similarly well as max-pooling, on RTE it even achieves higher performance than max-pooling. In the end, max-pooling seems to be a better choice as mean-pooling as it works better for most GLUE tasks.
Model GLUE avg.
RoBERTa Base 82.62
$\text{Max}_{\text{Seq}}$+MHA 82.6
$\Delta$ -0.02
Table 5: Average performance across all GLUE tasks. MaxPoolBERT shows a gain over RoBERTa base.
@@
Additionally we apply the max-pooling and MHA combination to a BERT variant, namely RoBERTa (Liu et al., 2019). The results are depicted in Table 6. For RoBERTa, our max-pooling + MHA variant shows improvements on 4 out of 9 tasks but on average both models perform equally (see Table 5). RoBERTa is an optimized version of BERT with a similar architecture, however the effects vary on this BERT variant. A significant improvement can only be observed on the CoLA dataset.
Model CoLA MRPC QNLI MNLI SST-2 STS-B RTE QQP WNLI
MCC Acc. F1 Acc. m $m$ Acc. Sp. Acc. Acc. F1 Acc.
[…]
RoBERTa Base 54.96 88.56 91.64 92.29 87.15 86.97 93.2 89.78 73.29 90.84 87.64 56.34
$\text{Max}_{\text{Seq}}$+MHA 57.43 87.58 91.01 92.15 86.99 87.14 93.43 90.49 70.4 90.93 87.85 55.87
$\Delta$ 2.47 -0.98 -0.63 -0.14 -0.16 0.17 0.23 0.71 -2.89 0.09 0.21 -0.47
Table 6: Performance on GLUE validation tasks for RoBERTa (average of 3 seeds).
@@ 6 Conclusion
We introduced MaxPoolBERT, a lightweight yet effective refinement of BERT’s classification pipeline that improves the representational quality of the [CLS] token for the BERT base model. Our method leverages max-pooling across layers and tokens, and introduces a multi-head attention layer that allows the [CLS] token to re-aggregate contextual information before classification. These modifications require no changes to pre-training and add minimal overhead to fine-tuning.
Empirical results on the GLUE benchmark demonstrate that MaxPoolBERT outperforms standard BERT base across most tasks, with especially strong improvements in low-resource settings. This suggests that BERT’s native use of the final-layer [CLS] embedding underutilizes available information and that small architectural additions can enhance generalization without sacrificing efficiency.
@@
  • Model size and generalization. Our work focuses on BERT base. We also examined one BERT variant but were not able to demonstrate an advantage of applying max-pooling + MHA for RoBERTa.
@@
In the future we aim to further investigate how to optimize the fine-tuning of small BERT models. While larger models often yield better performance, smaller models are crucial in real-time or resource-constrained environments. The BERT training paradigm following pre-training and fine-tuning has been predominant for several years and is widely used, so it is important to study whether further improvements can be made through small changes to this learning paradigm.
Acknowledgments
This research has been funded/supported by the Federal Ministry of Education and Research of Germany and the state of North Rhine-Westphalia as part of the Lamarr Institute for Machine Learning and Artificial Intelligence.
@@ A.1 Ablations
We conduct ablation studies to test different modifications of our architectures. We describe all experiments and report their results in the following.
@@ A.1.1 Choice of $k$
A.1.2 Time Difference
To evaluate differences in fine-tuning and inference time, we measured the time to fine-tune both standard BERT and MaxPoolBERT for four epochs on the MRPC dataset on a single A100 GPU. We also measured the inference time on the MRPC validation set for both model variants. Fine-tuning BERT took 289.298 seconds (approx. 72.324 seconds per epoch), inference on the validation set took 2.9935 seconds. In contrast, fine-tuning MaxPoolBERT on the MRPC dataset takes: 294.504 seconds (approx. 73.626 seconds per epoch). Inference time on the validation set is 3.0284 seconds. That is a difference of approximately 1.3 seconds per epoch of fine-tuning and 0.035 seconds difference for inference, which is neglectable.